OpenStudy (anonymous):

Will Medal and fan... Just need an explanation please... What is the solution of the equation? 3 ^5sqrt(x+2)^3 +3=27

OpenStudy (anonymous):

\[3\sqrt[5]{x+2}^{3}+3=27\]

OpenStudy (anonymous):

\[x= \frac{ 4 }{ 9\times 3^{2/3} } - 2\] = -1,7863

OpenStudy (anonymous):

\[3\sqrt[5]{x+2}^3 + 3 = 27\]\[3\sqrt[5]{x+2}^3= 24\]\[\sqrt[5]{x+2}^3 =8\]\[\sqrt[5]{x+2} = 2\]\[x+2 = 32\]\[x=30\]